OpenStudy (yb1996):

use linear approximation to estimate cos(61 degrees)

OpenStudy (freckles):

So we need to approximate \[\cos(\frac{61 \cdot \pi}{180}) \\ \text{ we know } \cos(\frac{60 \pi}{180})=\cos(\frac{\pi}{3})=\frac{1}{2} \text{ so find the tangent line to } \\ f(x)=\cos(x) \text{ at } x=\frac{\pi}{3}\]

OpenStudy (yb1996):

so the answer is \[\frac{ 1 }{ 2 }+\frac{ \sqrt{3} }{ 2 }(\pi/180)\]

OpenStudy (yb1996):

or \[\frac{ 1 }{ 2 }+\frac{ \pi \sqrt{3} }{ 360 }\]

OpenStudy (freckles):

I think you should have that -sin(pi/3) is -sqrt(3)/2 and not sqrt(3)/2

OpenStudy (yb1996):

oh ok, that makes sense

OpenStudy (yb1996):

Thanks for the help!

OpenStudy (freckles):

np